OpenStudy (anonymous):

Can someone explain the steps to finding the points for an equation on a graph? For example, I need to graph y=2x, how would i find the points for it?

OpenStudy (anonymous):

Choose 2 points step 1 x=0, then y=2*0=0-> point 1 (0,0) step 2 x =1, then y=2*1 =2 -> point 2 (1,2)

OpenStudy (anonymous):

How about y=1/2x Thats y equals 1 over 2 (one half) x